# Simultaneous analysis of approximate leave-one-out cross-validation and mean-field inference

### v1

@@ Abstract.
Approximate Leave-One-Out Cross-Validation (ALO-CV) is a method that has been proposed to estimate the generalization error of a regularized estimator in the high-dimensional regime where dimension and sample size are of the same order, the so called “proportional regime”. A new analysis is developed to derive the consistency of ALO-CV for non-differentiable regularizer under Gaussian covariates and strong-convexity of the regularizer. Using a conditioning argument, the difference between the ALO-CV weights and their counterparts in mean-field inference is shown to be small. Combined with upper bounds between the mean-field inference estimate and the leave-one-out quantity, this provides a proof that ALO-CV approximates the leave-one-out quantity as well up to negligible error terms. Linear models with square loss, robust linear regression and single-index models are explicitly treated.
@@ 1. Introduction to ALO-CV and mean-field inference
for any candidate vector $b\in\mathbb{R}^{p}$, where $(x_{new},y_{new})$ is an independent copy of $(x_{i},y_{i})$. The goal of the present paper is to analyse simultaneously two statistical methods to estimate the generalization error of an estimate
@@
for some constant $\delta\in(0,1)$. This is commonly referred to as the proportional regime, of proportional asymptotics, which has been the focus of much recent research in high-dimensional statistics and optimization. We refer to [24] for an introduction to the topic. Let us survey two methods that have been proposed to estimate (1.2) in the regime (1.3).
@@ 1.1. Approximate leave-one-out cross-validation
Error bounds of the estimation error of (1.2) by the leave-one-out estimate (1.5) and consistency are obtained in [34, 35]. However, the leave-one-out estimate (1.5) requires to solve the minimization problem $n$ times. ALO-CV attempts to approximate (1.5) while solving the minimization problem only once on the full dataset. To do so, the ALO-CV estimate [25] is constructed by performing one Newton step, starting from $\hat{b}$, to minimize (1.4). The resulting approximation of $b^{i}$ is
$\tilde{b}_{\textsc{newton}}^{i}=\hat{b}+\Bigl{(}\sum_{k\in[n]\setminus\{i\}}x_{k}L_{y_{k}}^{\prime\prime}(x_{k}^{T}\hat{b})+\nabla^{2}R(\hat{b})\Bigr{)}^{-1}x_{i}L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})$ (1.6)
which is well-defined as long as $L_{y_{k}}$ is twice differentiable with second derivative $L_{y_{k}}^{\prime\prime}$ and the Hessian $\nabla^{2}R$ exists and is positive-definite at $\hat{b}$. Replacing $b^{i}$ by $\tilde{b}_{\textsc{newton}}^{i}$ in the leave-one-out estimate (1.5), and using The Woodbury-Sherman-Morrison matrix identity to obtain
@@
Each step above is detailed in [25, 1]. This construction has been extended to some regularizer that are not differentiable such as $R(b)=\lambda\|b\|_{1}$ [33, 32, 1, 25, §2.2].
ALO-CV is appealing because its construction is simply based on the leave-one-out estimate and the above Newton step, with no assumption on the underlying data-generating process. It is computationally efficient compared to the leave-one-out estimate (1.5). It does not suffer from the sample size bias seen when using 5- or 10-fold cross-validation to approximate (1.2) [34, 25, Figure 1]: Estimating (1.2) by 5-fold cross-validation produces approximations the generalization error of an estimator trained on $4/5$ of the data, which is not the same as the generalization error of (1.1) trained on the full dataset.
Proving the success ALO-CV for estimating (1.5) or (1.2), for instance consistency, was initially obtained for Gaussian designs under differentiability and smoothness conditions on the loss functions $L_{y_{i}}$ and the penalty $R$ [25, 34]. The differentiability and smoothness conditions were recently relaxed to allow for the elastic-net penalty under stability conditions on the supports $S^{i}=\{j\in[p]:b^{i}_{j}\neq 0\}$ of (1.4) [1, Theorem 3.2], and such stability condition is shown to hold for the square loss $L_{y_{i}}(t)=(y_{i}-t)^{2}/2$ and the elastic-net penalty under isotropic Gaussian $x_{i}$ [1, §4]. The work [1] leverages these support stability results and smoothing argument to prove consistency of ALO-CV.
This paper presents a completely different argument, without relying on smoothing arguments or support stability conditions. The main contribution of the paper is a simple argument to prove the consistency of ALO-CV for non-smooth regularizer under Gaussian design, by directly relating the ALO-CV estimate (1.7) to results obtained in mean-field asymptotics that we now describe.
@@ 1.2. Mean-field asymptotics
The last decade has seen the development of mean-field asymptotics, which aims to study the asymptotic behavior of high-dimensional estimators such as (1.1) in the proportional regime (1.3), see for instance [17, 2, 19, 29, 16, 18, 30, 23, 14, 22] and [24] for an introduction to the topic. The picture that emerges is that the behavior of the empirical distribution of $\hat{b}$ or of $(x_{i}^{T}\hat{b})_{i\in[n]}$ can be precisely described by a few deterministic scalars that are the solution of a system of nonlinear equations. The precise definition of these deterministic quantities is not relevant for the purpose of the present paper and the aforementioned papers provide numerous examples. For the purpose of the present paper, let us mention that under regularity conditions on some test function $g$ of interest, if $x_{i}\sim N(0,I_{p})$ and $y_{i}\mid x_{i}$ follows a single-index model in the sense that $y_{i}=F(x_{i}^{T}w,E_{i})$ for some deterministic $F(\cdot,\cdot)$ and $w\in\mathbb{R}^{p}$ with $\|w\|_{2}=1$ as well as some external randomness $E_{i}$ independent of $x_{i}$, these results typically provide under Gaussian design and suitable assumptions
@@ 1.3. Mean-field inference
One family of results in mean-field asymptotics focuses on inference regarding the components $\beta_{j}^{*}$ of a ground-truth $\beta^{*}$ in regression models, or of the components $w_{j}$ of a ground-truth $w$ in single-index models. This is referred to as de-biasing, where the initial estimate $\hat{b}_{j}$ provided by the $j$-th component of (1.1) requires a bias correction after which the de-biased estimate is approximately normal and centered at $\beta_{j}^{*}$ in regression [8, 9, 14], or at a shrunk version of or $w_{j}$ in single index models [3, Section 4.1].
Another family of results useful for inference lets us construct estimate of the generalization error (1.2). The principle here is similar to de-biasing: for a given coordinate $i\in[n]$, start with the initial value $x_{i}^{T}\hat{b}$ and find a de-biasing correction that makes $x_{i}^{T}\hat{b}+\textsf{correction}$ approximately normal, see for instance [3, Theorem 4.3]. Let us state formally one version of such results that will be useful for the present paper.
@@
Furthermore, $e_{i}\in\mathbb{R}^{n}$ is the $i$-th canonical basis vector in $\mathbb{R}^{n}$, and $e_{j}\in\mathbb{R}^{p}$ the $j$-th canonical basis vector in $\mathbb{R}^{p}$. We say that the regularizer is $(\mu n,\Sigma)$-strongly convex if $b\mapsto R(b)-\mu n\|\Sigma^{1/2}b\|_{2}^{2}/2$ is convex, or equivalently
@@
Consider iid $(x_{i},y_{i})_{i\in[n]}$ with $x_{i}\sim N(0,\Sigma)$, and $w\in\mathbb{R}^{p}$ with $\mathbb{E}[(w^{T}x_{i})^{2}]\in\{0,1\}$ such that $(I_{p}-\Sigma ww^{T})x_{i}$ is independent of $(x_{i}^{T}w,y_{i})$. Assume that for all value of $y\in\mathcal{Y}$, the loss $L_{y}(\cdot)$ is differentiable and 1-Lipschitz, with $L_{y}^{\prime}(\cdot)$ being 1-Lipschitz. Assume that the regularizer $R$ is $(n\mu,\Sigma)$-strongly convex in the sense of (1.10) and that $R$ is minimized at 0. Then for fixed values of $(y_{i})_{i\in[n]}$ the mapping
@@
for some constant depending on $\mu,\delta$ only provided that $p/n$ satisfies (1.3). If additionally $g$ satisfies $\sup_{y\in\mathcal{Y}}|g(x,y)-g(x,y^{\prime})|\leq|x-x^{\prime}|(1+|x|+|x^{\prime}|)$ then
@@
Since $x_{i}\sim N(0,I_{p})$ is independent of $b^{i}$, (1.12) shows that $x_{i}^{T}b^{i}$ plus the additive correction $\operatorname{tr}[\Sigma\hat{A}]L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})$ is approximately normal; informally
@@
and $\|b^{i}\|_{2}$ in the right-hand side can be replaced by $\|\hat{b}\|_{2}$ up to a small error term by (1.14) below. The statement of Proposition 1.1 and its proof are variants of Theorem 4.3 in [3] with a minor variation on how the normal random variable is constructed. Here the standard normal random variable $x_{i}^{T}b^{i}/\|b^{i}\|_{2}$ is obtained by leaving one observation out, in [3, Theorem 4.3] it is obtained by replacing the $i$-th row of $X\in\mathbb{R}^{n\times p}$ by an independent copy. Since the results are not exactly, let us provide a short proof.
@@
Let us prove the case $\Sigma=I_{p}$. The case $\Sigma\neq I_{p}$ can be deduced from the case $\Sigma=I_{p}$ by a change of variable as explained in [5, Appendix B] or [3, Appendix C.2].
The derivative formula, the existence of the matrix $\hat{A}$ and the upper bound on $\|\hat{A}\|_{op}$ are proved in [3, Proposition 3.1] (or in [5] in linear regression). Let $P=I_{p}-ww^{T}$. Applying the consequence [7, (2.6)] of the Second Order Stein formula to the Gaussian vector $Px_{i}$ conditionally on $(x_{i}^{T}w,y_{i})$ gives
@@
For the first term, we use [1, Lemma 3 item 4][18, Appendices 2-3]:
$n\mu\|\hat{b}-b^{i}\|_{2}\leq\|x_{i}\|_{2}|L_{i}(x_{i}^{T}\hat{b})|.$ (1.14)
A short proof is as follows: write the KKT conditions $\sum_{l=1}^{n}x_{l}L_{y_{l}}^{\prime}(x_{l}^{T}\hat{b})\in-\partial R(\hat{b})$ for $\hat{b}$, the KKT conditions $\sum_{l\neq i}x_{l}L_{y_{l}}^{\prime}(x_{l}^{T}b^{i})\in-\partial R(b^{i})$ for $b^{i}$, and use the strong convexity property $(\hat{b}-b^{i})(\partial R(\hat{b})-\partial R(b^{i}))\geq n\mu\|\hat{b}-b^{i}\|_{2}^{2}$:
$n\mu\|\hat{b}-b^{i}\|_{2}^{2}$ $\leq(\hat{b}-b^{i})(\partial R(\hat{b})-\partial R(b^{i}))$
$=(\hat{b}-b^{i})^{T}x_{i}L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})-\sum_{l\neq i}\Bigl{(}x_{l}^{T}\hat{b}^{T}-x_{l}^{T}b^{i}\Bigr{)}\Bigl{(}L_{y_{l}}^{\prime}(x_{l}^{T}\hat{b})-L_{y_{l}}^{\prime}(x_{l}^{T}b^{i})\Bigr{)}.$
See [4, Section 4.2] where similar arguments are used to establish smoothness properties of $\hat{b}$. Since the rightmost sum is positive by convexity of $L_{y_{l}}$, (1.14) holds by the Cauchy-Schwarz inequality. Following the idea in [4, (C.3)], writing the strong convexity property of $R$ at $\hat{b}$ and 00 for the element of $\partial R(\hat{b})$ given by the KKT conditions further gives $n\mu\|\hat{b}-0\|_{2}^{2}\leq(\hat{b}-0)(\partial R(\hat{b})-\partial R(0))=\hat{b}^{T}\sum_{i=1}^{n}x_{i}L_{y_{i}}^{\prime}(x_{i}^{T}\hat{)}$ so that for instance
$n\mu\|\Sigma^{1/2}\hat{b}\|_{2}\leq\|X\Sigma^{-1/2}\|_{op}\sqrt{n}\max_{l\in[n]}|L_{y_{l}}(x_{l}^{T}\hat{b})|,$ (1.15)
by the same argument $n\mu\|\Sigma^{1/2}b^{i}\|_{2}\leq\|X\Sigma^{-1/2}\|_{op}\sqrt{n}\max_{l\in[n]\setminus\{i\}}|L_{y_{l}}(x_{l}^{T}b^{i})|.$ (1.16)
@@
$\mathbb{E}\Bigl{[}\sum_{i=1}^{n}B_{i}^{2}\Bigr{]}^{\frac{1}{2}}+\mathbb{E}\Bigl{[}\sum_{i=1}^{n}(U_{i}w^{T}(\hat{b}-b^{i}))^{2}\Bigr{]}^{\frac{1}{2}}+\mathbb{E}\Bigl{[}\|\hat{A}\|_{op}^{2}\sum_{i=1}^{n}L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})^{2}\Bigr{]}^{\frac{1}{2}}+\mathbb{E}\Bigl{[}\|\hat{b}^{T}P\hat{A}XD\|_{F}^{2}\Bigr{]}^{\frac{1}{2}}.$
Combining these upper bounds with (1.11), (1.14), (1.15), inequalities $|L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})|\leq 1$ and $|L_{y_{i}}^{\prime\prime}(x_{i}^{T}\hat{b})|\leq 1$ (so that $\|D\|_{op}\leq 1$) by assumption on $L_{y_{i}}$, the fact that $\mathbb{E}[\|X\|_{op}^{2}]\leq(\sqrt{n}+\sqrt{p})^{2}+1$ thanks to [15, Theorem II.13], equality $\mathbb{E}[\sum_{i=1}^{n}U_{i}^{2}\|x_{i}\|_{2}^{2}]\leq n(p-1)+3n$ since $x_{i}\sim N(0,I_{p})$ and $U_{i}=x_{i}^{T}w$, the claim (1.12) is proved. For (1.13), by assumption on the test function $g$ the left-hand side is bounded from above by $\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}|C_{i}|(1+|x_{i}^{T}b^{i}|+|C_{i}+x_{i}^{T}b^{i}|)$ so that the Cauchy-Schwarz inequality combined with (1.12) and (1.16) gives (1.13). ∎

### v2

@@ Abstract.
Approximate Leave-One-Out Cross-Validation (ALO-CV) is a method that has been proposed to estimate the generalization error of a regularized estimator in the high-dimensional regime where dimension and sample size are of the same order, the so-called “proportional regime”. A new analysis is developed to derive the consistency of ALO-CV for non-differentiable regularizers under Gaussian covariates and strong convexity. Using a conditioning argument, the difference between the ALO-CV weights and their counterparts in mean-field inference is shown to be small. Combined with upper bounds between the mean-field inference estimate and the leave-one-out quantity, this provides a proof that ALO-CV approximates the leave-one-out quantity up to negligible error terms. Linear models with square loss, robust linear regression and single-index models are explicitly treated.
@@ 1. Introduction to ALO-CV and mean-field inference
for any candidate vector $b\in\mathbb{R}^{p}$, where $(x_{new},y_{new})$ is an independent copy of $(x_{i},y_{i})$. The goal of the present paper is to analyze simultaneously two statistical methods to estimate the generalization error of an estimate
@@
for some constant $\delta\in(0,1)$. This is commonly referred to as the proportional regime, or proportional asymptotics, which has been the focus of much recent research in high-dimensional statistics and optimization. We refer to [24] for an introduction to the topic. Let us survey two methods that have been proposed to estimate (1.2) in the regime (1.3).
@@ 1.1. Approximate leave-one-out cross-validation
Error bounds of the estimation error of (1.2) by the leave-one-out estimate (1.5) and consistency are obtained in [34, 35]. However, the leave-one-out estimate (1.5) requires one to solve the minimization problem $n$ times. ALO-CV attempts to approximate (1.5) while solving the minimization problem only once on the full dataset. To do so, the ALO-CV estimate [25] is constructed by performing one Newton step, starting from $\hat{b}$, to minimize (1.4). The resulting approximation of $b^{i}$ is
[…]
which is well-defined as long as $L_{y_{k}}$ is twice differentiable with second derivative $L_{y_{k}}^{\prime\prime}$ and the Hessian $\nabla^{2}R$ exists and is positive-definite at $\hat{b}$. Replacing $b^{i}$ by $\tilde{b}_{\textsc{newton}}^{i}$ in the leave-one-out estimate (1.5), and using the Sherman-Morrison-Woodbury matrix identity to obtain
@@
Each step above is detailed in [25, 1]. This construction has been extended to some regularizers that are not differentiable such as $R(b)=\lambda\|b\|_{1}$ [33, 32, 1, 25, §2.2].
ALO-CV is appealing because its construction is simply based on the leave-one-out estimate and the above Newton step, with no assumption on the underlying data-generating process. It is computationally efficient compared to the leave-one-out estimate (1.5). It does not suffer from the sample size bias seen when using 5- or 10-fold cross-validation to approximate (1.2) [34, 25, Figure 1]: Estimating (1.2) by 5-fold cross-validation produces approximations of the generalization error of an estimator trained on $4/5$ of the data, which is not the same as the generalization error of (1.1) trained on the full dataset.
Proving the success of ALO-CV for estimating (1.5) or (1.2), for instance consistency, was initially obtained for Gaussian designs under differentiability and smoothness conditions on the loss functions $L_{y_{i}}$ and the penalty $R$ [25, 34]. The differentiability and smoothness conditions were recently relaxed to allow for the elastic-net penalty under stability conditions on the supports $S^{i}=\{j\in[p]:b^{i}_{j}\neq 0\}$ of (1.4) [1, Theorem 3.2], and such stability condition is shown to hold for the square loss $L_{y_{i}}(t)=(y_{i}-t)^{2}/2$ and the elastic-net penalty under isotropic Gaussian $x_{i}$ [1, §4]. The work [1] leverages these support stability results and smoothing arguments to prove consistency of ALO-CV.
This paper presents a completely different argument, without relying on smoothing arguments or support stability conditions. The main contribution of the paper is a simple argument to prove the consistency of ALO-CV for non-smooth regularizers under Gaussian design, by directly relating the ALO-CV estimate (1.7) to results obtained in mean-field asymptotics that we now describe.
@@ 1.2. Mean-field asymptotics
The last decade has seen the development of mean-field asymptotics, which aims to study the asymptotic behavior of high-dimensional estimators such as (1.1) in the proportional regime (1.3), see for instance [17, 2, 19, 29, 16, 18, 30, 23, 14, 22] and [24] for an introduction to the topic. The picture that emerges is that the behavior of the empirical distribution of $\hat{b}$ or of $(x_{i}^{T}\hat{b})_{i\in[n]}$ can be precisely described by a few deterministic scalars that are the solution of a system of nonlinear equations. The precise definition of these deterministic quantities is not relevant for the purpose of the present paper and the aforementioned papers provide numerous examples. For the purpose of the present paper, let us only describe one convergence result that is typical of mean-field asymptotics: If $x_{i}\sim N(0,I_{p})$, if the test function $g$ satisfies some regularity conditions, and if $y_{i}\mid x_{i}$ follows a single-index model in the sense that $y_{i}=F(x_{i}^{T}w,E_{i})$ for some deterministic $F(\cdot,\cdot)$ and $w\in\mathbb{R}^{p}$ with $\|w\|_{2}=1$ as well as some external randomness $E_{i}$ independent of $x_{i}$, then
@@ 1.3. Mean-field inference
One family of results in mean-field asymptotics focuses on inference regarding the components $\beta_{j}^{*}$ of a ground-truth $\beta^{*}$ in regression models, or of the components $w_{j}$ of a ground-truth $w$ in single-index models. This is referred to as de-biasing, where the initial estimate $\hat{b}_{j}$ provided by the $j$-th component of (1.1) requires a bias correction after which the de-biased estimate is approximately normal and centered at $\beta_{j}^{*}$ in regression [8, 9, 14], or at a shrunk version of $w_{j}$ in single index models [4, Section 4.1].
Another family of results useful for inference lets us construct an estimate of the generalization error (1.2). The principle here is similar to de-biasing: for a given coordinate $i\in[n]$, start with the initial value $x_{i}^{T}\hat{b}$ and find a de-biasing correction that makes $x_{i}^{T}\hat{b}+\textsf{correction}$ approximately normal, see for instance [4, Theorem 4.3]. Let us state formally one version of such results that will be useful for the present paper.
@@
Furthermore, $e_{i}\in\mathbb{R}^{n}$ is the $i$-th canonical basis vector in $\mathbb{R}^{n}$, and $e_{j}\in\mathbb{R}^{p}$ the $j$-th canonical basis vector in $\mathbb{R}^{p}$. We say that the regularizer is $(n\mu,\Sigma)$-strongly convex if $b\mapsto R(b)-\mu n\|\Sigma^{1/2}b\|_{2}^{2}/2$ is convex, or equivalently
@@
Consider iid $(x_{i},y_{i})_{i\in[n]}$ with $x_{i}\sim N(0,\Sigma)$, and $w\in\mathbb{R}^{p}$ with $\mathbb{E}[(w^{T}x_{i})^{2}]\in\{0,1\}$ such that $(I_{p}-\Sigma ww^{T})x_{i}$ is independent of $(x_{i}^{T}w,y_{i})$. Assume that for all values of $y\in\mathcal{Y}$, the loss $L_{y}(\cdot)$ is differentiable and 1-Lipschitz, with $L_{y}^{\prime}(\cdot)$ being 1-Lipschitz. Assume that the regularizer $R$ is $(n\mu,\Sigma)$-strongly convex in the sense of (1.10) and that $R$ is minimized at 0. Then for fixed values of $(y_{i})_{i\in[n]}$ the mapping
@@
for some constant depending on $\mu,\delta$ only provided that $p/n$ satisfies (1.3). If additionally $g$ satisfies $\sup_{y\in\mathcal{Y}}|g(x,y)-g(x,y)|\leq|x-x^{\prime}|(1+|x|+|x^{\prime}|)$ then
@@
[…]
@@
and $\|b^{i}\|_{2}$ in the right-hand side can be replaced by $\|\hat{b}\|_{2}$ up to a small error term by (1.14) below. The statement of Proposition 1.1 and its proof are variants of Theorem 4.3 in [4] with a minor variation on how the normal random variable is constructed. Here the standard normal random variable $x_{i}^{T}b^{i}/\|b^{i}\|_{2}$ is obtained by leaving one observation out, in [4, Theorem 4.3] it is obtained by replacing the $i$-th row of $X\in\mathbb{R}^{n\times p}$ by an independent copy. Since Proposition 1.1 does not appear in the literature in the form stated, let us provide a short proof.
@@
Let us prove the case $\Sigma=I_{p}$. The case $\Sigma\neq I_{p}$ can be deduced from the case $\Sigma=I_{p}$ by a change of variable as explained in [5, Appendix B] or [4, Appendix C.2].
The derivative formula, the existence of the matrix $\hat{A}$ and the upper bound on $\|\hat{A}\|_{op}$ are proved in [4, Proposition 3.1] (or in [5] in linear regression). Let $P=I_{p}-ww^{T}$. Applying the consequence [7, (2.6)] of the Second Order Stein formula to the Gaussian vector $Px_{i}$ conditionally on $(x_{i}^{T}w,y_{i})$ gives
@@
For the first term $\|P(\hat{b}-b^{i})\|^{2}$, we use [1, Lemma 3 item 4][18, Appendices 2-3]:
[…]
$n\mu\|\hat{b}-b^{i}\|_{2}^{2}$ $\leq(\hat{b}-b^{i})^{T}(\partial R(\hat{b})-\partial R(b^{i}))$
$=(\hat{b}-b^{i})^{T}\Bigl(-x_{i}L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})\Bigr)-\sum_{l\neq i}\Bigl(x_{l}^{T}\hat{b}-x_{l}^{T}b^{i}\Bigr)\Bigl(L_{y_{l}}^{\prime}(x_{l}^{T}\hat{b})-L_{y_{l}}^{\prime}(x_{l}^{T}b^{i})\Bigr).$
[…]
by the same argument $n\mu\|\Sigma^{1/2}b^{i}\|_{2}\leq\|X\Sigma^{-1/2}\|_{op}\sqrt{n}\max_{l\in[n]\setminus\{i\}}|L_{y_{l}}^{\prime}(x_{l}^{T}b^{i})|.$ (1.16)
@@
$\mathbb{E}\Bigl[\sum_{i=1}^{n}B_{i}^{2}\Bigr]^{\frac{1}{2}}+\mathbb{E}\Bigl[\sum_{i=1}^{n}(U_{i}w^{T}(\hat{b}-b^{i}))^{2}\Bigr]^{\frac{1}{2}}+\mathbb{E}\Bigl[\|\hat{A}\|_{op}^{2}\sum_{i=1}^{n}L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})^{2}\Bigr]^{\frac{1}{2}}+\mathbb{E}\Bigl[\|\hat{b}^{T}P\hat{A}X^{T}D\|_{F}^{2}\Bigr]^{\frac{1}{2}}.$
The claim (1.12) is proved by combining these upper bounds with (1.11), (1.14), (1.15), inequalities $|L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})|\leq 1$ and $|L_{y_{i}}^{\prime\prime}(x_{i}^{T}\hat{b})|\leq 1$ (so that $\|D\|_{op}\leq 1$) by assumption on $L_{y_{i}}$, the fact that $\mathbb{E}[\|X\|_{op}^{2}]\leq(\sqrt{n}+\sqrt{p})^{2}+1$ thanks to [15, Theorem II.13], and the inequality $\mathbb{E}[\sum_{i=1}^{n}U_{i}^{2}\|x_{i}\|_{2}^{2}]\leq n(p-1)+3n$ since $x_{i}\sim N(0,I_{p})$ and $U_{i}=x_{i}^{T}w$. For (1.13), by assumption on the test function $g$ the left-hand side is bounded from above by $\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}|C_{i}|(1+|x_{i}^{T}b^{i}|+|C_{i}+x_{i}^{T}b^{i}|)$ so that the Cauchy-Schwarz inequality combined with (1.12) and (1.16) gives (1.13). ∎